OpenStudy (anonymous):

trig derivative question. I'm trying to find the derivative for the following function but I'm uncertain which trig identity to choose. \[f(x)=\cos^{2}(2x)-5\]

OpenStudy (anonymous):

There's the double angle identity \[\cos(2a)=\cos^{2}(a)-\sin^{2}(a)\] and \[\cos^{2}(a)=\frac{1}{2}(1+\cos(2a))\]